OpenStudy (khally92):

CALCULUS 3. Having Difficulty setting up. Use a double integral to find the exact volume of the wedge cut from the cylinder x^2+y^2=4 by the planes z=0 and z=2-y Thank you.

OpenStudy (inkyvoyd):

draw the region first. ALWAYS draw the region if you can You can either choose to draw the domain of integration, or draw the entire volume. Up to you.

OpenStudy (khally92):

That where the problme is.

OpenStudy (inkyvoyd):

What's the issue?

OpenStudy (khally92):

I know Z=0 is the XY plane

OpenStudy (inkyvoyd):

and?

OpenStudy (khally92):

Z=2-y is a straight line.

OpenStudy (inkyvoyd):

is it a straight line?

OpenStudy (inkyvoyd):

what values can x take?

OpenStudy (khally92):

you mean for the cylinder.

OpenStudy (inkyvoyd):

No. For z=2-y

OpenStudy (khally92):

but there is no x in the equation, x=0

OpenStudy (inkyvoyd):

oh? why can't x be equal to 1? Is the equation true if x=0? if x=1? how about x=2?

OpenStudy (khally92):

The equation is true for x=0, because the line is on the yz plane. x cannot take any value.

OpenStudy (khally92):

??

OpenStudy (inkyvoyd):

Is there anything in the equation saying that x has to be equal to 0?

OpenStudy (khally92):

Nope.

OpenStudy (inkyvoyd):

Then how do you know?

OpenStudy (khally92):

Why are we talking about x here.

OpenStudy (khally92):

we don't have x.

OpenStudy (inkyvoyd):

Because, you cannot do the problem without your domain of integration. you assume that z=2-y graphs a straight line, but how are you supposed to do the problem in R^3 when you have a line, a plane, and a cylinder? surely one of the assumptions you have made is wrong.

OpenStudy (inkyvoyd):

let's look at z=0 what does that graph? what values can y, and x take on?

OpenStudy (khally92):

Z=0 y=2

OpenStudy (inkyvoyd):

no, I'm talking about what you said was the xy plane. z=0

OpenStudy (khally92):

its a plane xy-plane

OpenStudy (inkyvoyd):

yes. What values can x take? what values can y take?

OpenStudy (inkyvoyd):

According to your logic for z=2-y, x=0 and y=0 too.

OpenStudy (khally92):

z=0 is the whole xy plane both x and y can have any value

OpenStudy (khally92):

well at least positive

OpenStudy (inkyvoyd):

then what is z=2-y?

OpenStudy (khally92):

on 2D or 3D

OpenStudy (inkyvoyd):

and are you sure x and y have to be positive? we are talking about a plane, not a quarter plane

OpenStudy (inkyvoyd):

What do you think the problem is referring to? 2D or 3D, you judge

OpenStudy (khally92):

I think everything here is in 3D?

OpenStudy (khally92):

but not for the domain

OpenStudy (inkyvoyd):

so what is z=2-y?

OpenStudy (khally92):

A Plane

OpenStudy (inkyvoyd):

so, what does the domain look like?

OpenStudy (khally92):

from 0 to the plane z=y-2

OpenStudy (inkyvoyd):

that is for z yes. what about x and y?

OpenStudy (khally92):

I am totally lost

OpenStudy (inkyvoyd):

Okay. So the first thing we always do is sketch the region, or the domain we integrate in, no? can you do this?

OpenStudy (khally92):

I cannot sketch the region, and cant figure the domain either.

OpenStudy (khally92):

when you say region you mean the cylinder?

OpenStudy (inkyvoyd):

No. We are doing a double integral here, so we don't actually integrate the cylinder with a density function of one. We integrate a CUT of the cylinder that is more or less of constant cross sectional area... take your guess which plane we cut the cylinder in

OpenStudy (khally92):

the cylinder is on the z axis and the plane z=2-y cut the cylinder ?

OpenStudy (inkyvoyd):

No. we actually integrate over the region x^2+y^2=4

OpenStudy (khally92):

Oh yeah I know that.

OpenStudy (khally92):

The headache is finding the domain?

OpenStudy (khally92):

which i cannot still figure out.

OpenStudy (inkyvoyd):

\(\Huge \int^?_?\int^?_??d?d?\) fill in what you know

OpenStudy (inkyvoyd):